If a card is drawn from an ordinary deck, find he probability of drawing each of the following spade or
face card, and face or red card?

Answer by rothauserc(4718)

In an ordinary deck there are 52 cards, 4 suits of 13 cards each and 12 face cards(3 face cards per suit)
:
Probability(P) of drawing a spade or face card = (13/52) +(12/52) -(3/52) = 22/52 = 11/26
:
Note We subtract the 3 spade face cards since they were already counted in 13/52
:
P of drawing a red card or face card = (26/52) +(12/52) -(6/52) = 32/52 = 8/13
: